OpenStudy (anonymous):

If f(x) is a linear function f(5) = 51 and f'(5) = 3, what is the value of f(-3) + f'(-3)?

OpenStudy (watchmath):

Let say \(f(x)=ax+b\). What is \(a\) and \(b\) in order to satisfy the requirement?

OpenStudy (anonymous):

I always get f(-3) - f'(-3) = 45 can you check that out?

OpenStudy (watchmath):

Isn't the question asking for f(-3)+f'(-3) ?

OpenStudy (anonymous):

oh sorry that's typo

OpenStudy (watchmath):

\(f'(x)=a\) for all \(x\). SInce \(f'(5)=3\) then \(a=3\). Now \(51=f(5)=3\cdot 5+b\). So \(b=36\). Now you can find \(f(-3)+f'(-3)\)

OpenStudy (anonymous):

I suddenly realized my mistake thanks =)